OpenStudy (anonymous):

Find the vertex, axis of symmetry, domain, and range of the graph of y = −3x2 − 3x + 4 Explain how you can determine the axis of symmetry, the domain, and range without graphing y = −3x2 − 3x + 4.

OpenStudy (anonymous):

use \(x=-\frac{b}{2a}\) for your axis of symmetry

OpenStudy (anonymous):

in your case \(y=-3x^2-3x+4\) you have \(a=-3,b=-3, -\frac{b}{2a}=-\frac{-3}{2\times (-3)}=-\frac{1}{2}\)

OpenStudy (anonymous):

is the axis of symmetry -1/2 ?

OpenStudy (anonymous):

yup well the axis of symmetry is not a number it is a line the equation for the line is \(x=-\frac{1}{2}\)

OpenStudy (anonymous):

equation for a vertical like is \(x=c\) for some constant \(c\)

OpenStudy (anonymous):

domain is easiest it is a polynomial, so domain is all real numbers

OpenStudy (anonymous):

so is the range y<= 19/4?

OpenStudy (anonymous):

for the range you need to see what you get when you replace \(x\) by \(-\frac{1}{2}\) since this is a parabola that opens down, the range will be from minus infinity to whatever you get for the \(y\) value at \(x=-\frac{1}{2}\)

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

thanks! can you help with part 2?? @satellite73